OpenStudy (anonymous):

Solve for x: |2x + 6| − 4 = 20

OpenStudy (anonymous):

Since it is an absolute value function, you must divide it into 2 parts, the first one is 2x+6 and the other one is -(2x+6). Solve the 2 equations right away and tell me your answer, i already got mine. :)

OpenStudy (anonymous):

These are the choices x = 9 and x = 11 x = −9 and x = 15 x = 9 and x = −15 No solution

OpenStudy (nurali):

|2x + 6| − 4 = 20 |2x + 6| =20+4 |2x + 6| =24 2x + 6=24 or 2x + 6=-24 2x=24-6 or 2x=-24-6 2x=18 or 2x=-30 x=18/2 or x=-30/2 x=9 or x=-15

OpenStudy (nurali):

Welcome to Openstudy.

OpenStudy (anonymous):

9 and -15